OpenStudy (anonymous):

i gotta show work 4 this: {x=2; y=4} 12/y + 9y/x how to solve?

OpenStudy (anonymous):

\[\frac{12}{y}+\frac{9y}{x}\]\[ x=2, y=4\]Simply substitute x,y with the values given\[\frac{12}{4}+\frac{9\times4}{2}=\]\[3+\frac{36}{2}=\]\[3+18=\]\[21\]